An elastic cord can be stretched to its elastic limit by a load of 2N. If a 35cm length of the cord is extended 0.6cm by a force of 0.5N what will be the length of the cord when the stretching force is 2.5N?


1
Expert's answer
2022-03-30T03:31:54-0400

"k=\frac F{\Delta x}=83.3~\frac Nm,"

"\Delta x_1=\frac {F_1}k=\frac {F_1}F\Delta x=2.4~cm,"

"y=x_0+\Delta x_1=37 .4~cm."